# CLOVA: A Closed-LOop Visual Assistant with Tool Usage and Update

### v1

@@ 7.4 Prompts in learning
In the learning phase, we use LLMs to infer answers for the VQA model, and then train the VQA model using the question and inferred answer. One example of prompts sent to LLMs for answer inferring is shown in Fig. 16, respectively.
Figure 16: Prompts of inferring answers for the VQA model.
@@ 8.2.3 Inference
The prompt ensemble process of the LOC tool also has two steps. (1) Given a visual concept and a query image, we select all the prompts having the same visual concept from the prompt pool as candidates. We then filter out the prompts by making predictions with each candidate prompt, if the prediction confidence is larger than 0.1, we will use the prompt for the query image, otherwise, we will remove this candidate prompt. (2) We use prompt ensemble (detailed in Section 3.4) to aggregate all the selected prompts for a query instance and contact the aa’gag’gaggrag’g’reaggregaggregaaggrega prompt with the input to produce a prediction.
@@ 8.4.2 Training
We update the CLIP through the deep soft prompt tuning [18]. We utilize the Adam optimizer with a learning rate set to $21e-$ 2. To obtain training data, CLOVA uses the concept to be learned to automatically retrieve $7$ images from Google Images as positive data through web scraping. The first image is used for validation, while the remaining ones are used for training. Furthermore, CLOVA derives additional concepts related to but different from the target concept through GPT and subsequently collected data associated with these concepts as negative samples. CLOVA then uses both the positive data and negative data for training. During the training phase, we introduce $100$ learnable prompts for each of the first three layers of the vision transformer to facilitate soft prompt learning. We conduct prompt training for $100$ steps per instance, followed by a validation step. If the accurate prediction is achieved on the validation data, CLOVA systematically stores features of crawled images and learned prompts in the prompt pool, otherwise, these learned prompts are discarded.
@@ 8.5 Update REPLACE tool
We use Stable Diffusion [48] as the REPLACE tool. In Stable Diffusion, the architecture comprises four key components: Sampler, Variational Autoencoder (VAE), UNet, and CLIPEmbedder. The Sampler and UNet focus on the actual image generation, the VAE provides a deep understanding of image content, and the CLIPEmbedder ensures the relevance and accuracy of the generated images in relation to the text inputs. We added learnable prompts to the text encoder of the CLIPEmbedder component. During the prompt tuning phase, we tune a prompt vector for training images and store it. The dimensionality of the prompt vector is $768$. The architecture of Stable Diffusion is shown in Fig. 21.
@@ 8.5.2 Inference
Similar to other visual models, inference for the Stable Diffusion model also contains two steps. Given query data, CLOVA first locates prompts corresponding to the concept in the prompt pool and then loads the prompt into the text encoder of the Stable Diffusion model. Based on the query and mask, we perform editing on the input image.
@@ 9.2 Evaluation on the online setting
CLOVA can be applied to a more practical online learning setting. In this case, CLOVA is evaluated in a dynamic data stream. If it makes a correct prediction on a task, only the inference phase is activated, and the task is tagged as a correct prediction; if it makes an incorrect prediction on a task, this task is tagged as a wrong prediction, and the reflection and learning phases are activated to update tools. After the data stream, we calculate the accuracy based on tagged predictions of all cases. We conduct experiments on the compositional VQA and multi-image reasoning tasks, where the GQA and NLVRv2 datasets are used. Results are shown in Tab. 7. Similar to the offline setting in Section 4.2, updating LLMs and VFMs both leads to improvements.
@@
+ Update visual models $50.2$ $53.0$ $57.8$
@@
+ Update visual models $61.6$ $62.6$ $67.4$

### v2

@@ 8.4 Prompts in learning
In the learning phase, we use LLMs to infer answers for the VQA tool, and then tune prompts of the VQA tool using the question and inferred answer. One example of prompts sent to LLMs for answer inferring is shown in Fig. 16, respectively.
Figure 16: Prompts of inferring answers for the VQA tool.
@@ 9.2.3 Inference
The prompt ensemble process of the LOC tool also has two steps. (1) Given a visual concept and a query image, we select all the prompts having the same visual concept from the prompt pool as candidates. We then filter out the prompts by making predictions with each candidate prompt, if the prediction confidence is larger than 0.1, we will use the prompt for the query image, otherwise, we will remove this candidate prompt. (2) We use prompt ensemble (detailed in Section 3.4) to aggregate all the selected prompts for a query instance and contact the prompt with the input to produce a prediction.
@@ 9.4.2 Training
We update the CLIP through the deep prompt tuning [20]. We utilize the Adam optimizer with a learning rate set to $21e-$ 2. To obtain training data, CLOVA uses the concept to be learned to automatically retrieve $7$ images from Google Images as positive data through web scraping. The first image is used for validation, while the remaining ones are used for training. Furthermore, CLOVA derives additional concepts related to but different from the target concept through GPT and subsequently collected data associated with these concepts as negative samples. CLOVA then uses both the positive data and negative data for training. During the training phase, we introduce $100$ learnable prompts for each of the first three layers of the vision transformer to facilitate prompt tuning. We conduct prompt tuning for $100$ steps per instance, followed by a validation step. If the accurate prediction is achieved on the validation data, CLOVA systematically stores features of crawled images and learned prompts in the prompt pool, otherwise, these learned prompts are discarded.
@@ 9.5 Update REPLACE tool
We use Stable Diffusion [57] as the REPLACE tool. In Stable Diffusion, the architecture comprises four key components: Sampler, Variational Autoencoder (VAE), UNet, and CLIPEmbedder. The Sampler and UNet focus on the actual image generation, the VAE provides a deep understanding of image content, and the CLIPEmbedder ensures the relevance and accuracy of the generated images in relation to the text inputs. We added learnable prompts to the text encoder of the CLIPEmbedder component. During the prompt tuning phase, we tune a prompt for training images and store it. The dimensionality of the prompt is $768$. The architecture of Stable Diffusion is shown in Fig. 21.
@@ 9.5.2 Inference
Similar to other visual tools, inference for the Stable Diffusion model also contains two steps. Given query data, CLOVA first locates prompts corresponding to the concept in the prompt pool and then loads the prompt into the text encoder of the Stable Diffusion model. Based on the query and mask, we perform editing on the input image.
@@ 10.2 Evaluation on the online setting
CLOVA can be applied to a more practical online learning setting. In this case, CLOVA is evaluated in a dynamic data stream. If it makes a correct prediction on a task, only the inference phase is activated, and the task is tagged as a correct prediction; if it makes an incorrect prediction on a task, this task is tagged as a wrong prediction, and the reflection and learning phases are activated to update tools. After the data stream, we calculate the accuracy based on tagged predictions of all cases. We conduct experiments on the compositional VQA and multi-image reasoning tasks, where the GQA and NLVRv2 datasets are used. Results are shown in Tab. 8. Similar to the offline setting in Section 4.2, updating LLMs and visual tools both leads to improvements.
@@
+ Update visual tools $50.2$ $53.0$ $57.8$
@@
+ Update visual tools $61.6$ $62.6$ $67.4$